# The universal multiplicity function: counting halos and voids

### v1

@@ Appendix D Measuring the moving barrier
To speed up the analysis, we start exploring the parameter space of the effective barrier of Eq. (16) performing a MCMC with the approximated analytical multiplicity function, Eq. (3). We do it for several values of the minimum void radius. For each minimum radius, we verify to be in the regime in which the analytical multiplicity function is an accurate approximation of the exact one corresponding to the effective barrier. To do so, we compare the best-fit analytical multiplicity function, Eq. (3), with the numerical one obtained with the same moving barrier, Eq. (5)–(6). According to the precision of the analytical approximation Eq. (3) and to the uncertainty of the VSF from ICs, we select a minimum radius in which the analytical approximation can be safely used, in a very conservative way, listed in the last column of Tab. 2. Then we select the parameter space region to be further explored as the region in which the logarithm of the posterior distribution is larger that the maximum value minus 5: $\theta$ s.t. $\log(p_{\rm max})-p(x,\theta)>5$, where $\theta=(\alpha,\beta,\gamma)$ and $x$ the measured VSF. We then use this region in the parameter space as the flat prior for the analysis with the numerical multiplicity function, Eq. (5)–(6). We repeat this procedure for all the void thresholds considered. It is worth noting that Eq. (3) is able to fit the VSF from simulation ICs even at small radii, but beyond the radius in which the analytical approximation breaks, this is no more representative of the first crossing solution of a generic moving barrier. This means that even if in principle we can use Eqs. (3), to fit the parameters of the moving barrier Eq. (16), at scales smaller than the radius at which the analytical approximation breaks, we lose the map between this function and the effective barrier. Since the aim of Sec. 4 is to find the effective moving barrier, we proceed to measure it using the numerical multiplicity function.
To perform the MCMC up to the smallest possible radius allowed by ICs resolution, we use the numerical solution of the first crossing problem, Eqs. (5)–(6), implemented with the 3-parameters barrier Eq. (16). To obtain a numerical multiplicity function with enough precision to be used in such MCMC, we use $10^{8}$ Monte Carlo realizations of Eq. (5) for each $\alpha$, $\beta$, and $\gamma$ values, computed with a binning radius size at least 4 times larger than the binning used for the VFS measured in simulations ICs. This guarantees a numerical precision that is, at worst, less than $1/10$ the Poissonian uncertainty of the measured VSF. Since this is computationally expensive for a MCMC analysis, we interpolate the numerical multiplicity function over a sample of $\alpha$, $\beta$, and $\gamma$ values. To select the range in a convenient way, we produce a large number of numerical multiplicity over a random sapling of $\alpha$, $\beta$, and $\gamma$, selected within the parameter space region delimited by the MCMC analysis involving the analytical multiplicity function, as described before.
At each MCMC step, we interpolate over this sample finding the closest tetrad in which the point of coordinates ($\alpha$, $\beta$, $\gamma$) is embedded (which in most of the cases does not correspond to the 4 closest sample points). For interpolating, we weight the numerical multiplicity function corresponding to each point using the barycenter coordinates. We verified that, for this specific application, this interpolation method is more accurate than the standard ones over a grid. Moreover, it is simple to increase the sample resolution if necessary, by producing more realizations at random positions within the same parameter space region, without worrying about the grid.
As a last consideration, it is worth noting that the moving barrier is really in modeling the Lagrangian VSF measured from simulation ICs. In particular, the five larger radius bins are always enough to constrain the 3-parameter barrier and to predict the VSF down to the smaller radii. However, using all the available radius bins, the uncertainty in the posterior decreases.
$\delta(z_{\rm IC})$ $R_{\rm min}(2000)$ $R_{\rm min}(500)$ $R^{\rm A}_{\rm min}$
-0.035 11 2.5 4
-0.023 17 5 7
-0.016 17 7 11
-0.011 19 9 19
-0.008 22 11 28
-0.0064 30 9 35
-0.005 39 12 45
@@
Table 2: Minimum radii for each threshold explored. The first column, $\delta(z_{\rm IC})$, lists the actual density contrast in the simulation ICs at $z_{\rm IC}=99$; the second and third columns, $R_{\rm min}(2000)$ and $R_{\rm min}(500)$, list the minimum radii used for the analysis with the numerical multiplicity function, which corresponds to $\sim 1.5-2.5$ times the Poissonian; the last column, $R^{\rm A}_{\rm min}$, lists the minimum radius used for the MCMC analysis with the analytical approximation, Eq. (3), chosen in a conservative way. All radii are in $h^{-1}{\rm Mpc}$ unit.
Appendix E Redshift dependence of the HMF
@@
Figure 9: Dependence of the halo formation threshold $\delta_{\rm h}$ and the moving barrier parameter parameters of Eq. (16) as a function of $\sigma_{8}(z)$. The upper axis shows the corresponding redshift. Orange violins show the posterior distribution obtained at each DEMNUni snapshot considered, $z=0$, 0.47, 1.05, 1.46, 2.05. Blue solid lines and dashed areas show the best-fit and 68% CL of the scaling relations described in the text.
@@
In this Section we discuss the redshift dependence of the Eulerian HMF, as described in Sec. 5, obtained using Eq. (17) as prior to fit the formation barrier $\delta_{\rm h}$. The redshift dependence of the theoretical HMF is better expressed as a function of $\sigma_{8}(z)=\sqrt{C(R,R)}$ at a given redshift, with $R=8\,h^{-1}{\rm Mpc}$. In particular, $\sigma_{8}(z)=\sigma_{8}(z=0)D_{0}(z)/D_{0}(z=0)$, where $D_{0}(z)$ is the linear growth factor of perturbations. In this way, the normalization effects of the Lagrangian space and any redshift dependence are automatically included. Nevertheless, Fig. 9 shows that all the parameters of the effective barrier, Eq. (16) and $\delta_{\rm h}$ can be consistently considered redshift independent. The orange violins shows the posterior of the parameter at the $\sigma_{8}(z)$ values corresponding to the redshift values explored, $z=0$, 0.47, 1.05, 1.46, 2.05. Blue lines show the best-fit value and the shaded area the 68% CL. The best-fit values are the following:
$\delta_{\rm h}$ $=2.424\pm^{0.076}_{0.071}$ $\alpha$ $=1.209\pm^{0.008}_{0.007}$ (E1)
@@
It is worth noting that redshift independent barrier parameters entail a self-similar HMF, i.e. an HMF that depends on $S=\langle\delta^{2}>$ (and $D$) values only. Note that the Lagrangian space is the initial density field linearly evolved up to the redshift of interest with the linear growth factor. This means that the linear formation threshold $\delta_{\rm h}$ does not depend on the redshift, but the linear growth factor used to evolve the Lagrangian space does. In computations, this can be considered alternatively by using the linear power spectrum of the redshift considered, or using the power spectrum at a reference redshift, e.g. $z=0$, and extrapolating the formation threshold at the redshift of interest: $\delta_{\rm h}\rightarrow\delta_{\rm h}D_{0}(z=0)/D_{0}(z)$  (SMT). According to the same argument, the effective moving barrier can be used either with the linear power spectrum at the redshift considered, or with the power spectrum at a reference redshift and extrapolating the barrier, which means $\alpha\rightarrow\alpha D_{0}(z=0)/D_{0}(z)$, where we considered $z=0$ as the reference redshift. Moreover, any parameter multiplied or divided by $\sigma$, or more generally by any quantity derived from the power spectrum, must be corrected for the growth factor. Therefore, also $\beta\rightarrow\beta D_{0}(z=0)/D_{0}(z)$.
The blue curve in Fig. 10 shows the relative difference between the Eulerian spherical mapping of the effective barrier approach implemented in this work (blue), Eqs. (3), (16), fitting the halo threshold from Eq. (17) as described in Sec. 5. Each panel shows a different redshift, corresponding to the DEMNUni snapshots considered. The error bars and the hatched area show ±$\pm 1\sigma$ interval, dashed area shows the ±$\pm 1\%$ uncertainty. The agreement is almost always within $1\sigma$. Simirarly to the case of fitting the moving barrier for the Lagrangian VSF, Appendix D, the model presented here is really robust in modeling the Eulerian HMF measured from simulation. As in the previous case, the posterior distributions of the considered parameter are stable from the larger scale explored down to the resolution limit of the simulation. For comparison, the best-fit of the SMT model (orange lines) and the best-fit of the 4-parameters Tinker et al. (2008) model are shown. It can be noticed that the agreement of our model with the HMF measured in simulation is almost always as good as the Tinker et al. (2008) one. The main differences are the clearer physical interpretation and the smaller parameter space of the model presented in this work. Moreover, our model converges much faster in the MCMC fit than the 4-parameters Tinker et al. (2008) model. Note that the multiplicity function of Eq. (3) can be also used as a versatile fitting function with as many fitting parameters as desired, which modify $B_{s}$ and $B^{\prime}_{S}$ quantities.
@@
Figure 10: Relative difference between theoretical HMFs and with respect to the HMF measured in the DEMNUni in unit of Poissonian uncertainty: Eulerian spherical mapping of the effective barrier approach implemented in this work (blue), Eqs. (3), (16), fitting the halo threshold from Eq. (17); SMT best-fit model (orange); Tinker et al. (2008) best-fit model (green). Each panel shows a different redshift, $z=0$, 0.47, 1.05, 1.46, 2.05. Errorbars and hatched area show ±$\pm 1\sigma$ interval, dashed area shows the ±$\pm 1\%$ uncertainty.

### v2

@@ Appendix E Measuring the moving barrier
To speed up the analysis, we start exploring the parameter space of the effective barrier of Eq. (4.1) performing a MCMC with the approximated analytical multiplicity function, Eq. (3.2). We consider a Gaussian likelihood function,
$\log\left[{\cal L}({\cal D}|\Theta)\right]=-\frac{1}{2}\sum_{i}\left[\left(n^{\cal D}_{i}-n^{T}_{i}(\Theta)\right)^{2}\Sigma^{-1}_{i}+\log(\Sigma_{i})\right],$ (E.2)
@@
where the theoretical model is the theoretical multiplicity function integrated in the radius bins considered $n^{T}_{i}(\Theta)=\int_{S(R_{i})}^{S(R_{i-1})}f(S){\rm d}S$ with $\Theta=\{\alpha,\beta,\gamma\}$; the data vector with elements $n^{\cal D}_{i}$, which is the measured multiplicity function in the various $i^{\rm th}$ bin, $n^{\cal D}_{i}=N_{i}4\pi(\langle R\rangle_{i}/L_{\rm box})^{3}/3$, where $N_{i}$ is the number of voids with radius between $R_{i-1}$ and $R_{i}$ and $\langle R\rangle_{i}=(R_{i-1}+R_{i})/2$; the variance $\Sigma_{i}$ is given by the Poissonian uncertainty of the measured multiplicity function, $\Sigma_{i}=N_{i}\left[4\pi(\langle R\rangle_{i}/L_{\rm box})^{3}/3\right]^{2}$. The prior is a flat distribution over the 3 parameters explored. Note that the radius bins explored in this analysis are thin, with a thickness ranging from $0.5$ to 1 $h^{-1}$pc. We repeat the analysis exploring several minimum void radius values. For each minimum radius, we verify to be in the regime in which the approximated analytical multiplicity function is an accurate approximation of the exact one corresponding to the effective barrier. To do so, we compare the best-fit approximated multiplicity function, Eq. (3.2), with the numerical one obtained with the same moving barrier, Eq. (3.4)–(3.5). According to the precision of the analytical approximation Eq. (3.2) and to the uncertainty of the VSF from ICs, we select a minimum radius in which the analytical approximation can be safely used, in a very conservative way, listed in the last column of Tab. 2. Then we select the parameter space region to be further explored as the region in which the logarithm of the posterior distribution is larger that the maximum value minus 5: $\theta$ s.t. $\log(p_{\rm max})-\log\big{(}p(x,\theta)\big{)}>5$, where $\theta=(\alpha,\beta,\gamma)$ and $x$ the measured VSF. We then use this region in the parameter space as the flat prior for the analysis with the numerical multiplicity function, Eq. (3.4)–(3.5). We repeat this procedure for all the void thresholds considered. It is worth noting that both Eq. (3.14) and Eq. (3.2) are able to fit the VSF from simulation ICs even at small radii, but beyond the radius in which the analytical approximation breaks, this is no more representative of the first crossing solution of a generic moving barrier. This means that even if in principle we can use Eqs. (3.2), to fit the parameters of the moving barrier Eq. (4.1), at scales smaller than the radius at which the analytical approximation breaks, we lose the map between this function and the effective barrier. Since the aim of Sec. 4 is to find the effective moving barrier, we proceed to measure it using the numerical multiplicity function.
$\delta(z_{\rm IC})$ $R_{\rm min}(2000)$ $N_{\rm v}(2000)$ $R_{\rm min}(500)$ $N_{\rm v}(500)$ $R^{\rm A}_{\rm min}$
-0.035 11 361 2.5 92965 4
-0.023 17 419 5 14900 7
-0.016 17 9586 7 6703 11
-0.011 19 18419 9 3601 19
-0.008 22 16437 11 2031 28
-0.0064 30 6556 9 3198 35
-0.005 39 3185 12 1386 45
@@
Table 2: Minimum radii for each threshold explored. The first column, $\delta(z_{\rm IC})$, lists the actual density contrast in the simulation ICs at $z_{\rm IC}=99$; the second and fourth columns, $R_{\rm min}(2000)$ and $R_{\rm min}(500)$, list the minimum radii used for the analysis with the numerical multiplicity function, which contains to $\sim 1.5-2.5$ times the number of particles associated to the Poissonian noise (see text); the third and fifth columns, $N_{\rm v}(2000)$ and $N_{\rm v}(500)$, list the number of voids in the simulation ICs with radius larger then the minimum one; the last column, $R^{\rm A}_{\rm min}$, lists the minimum radius used for the MCMC analysis with the analytical approximation, Eq. (3.2), chosen in a conservative way. All radii are in $h^{-1}{\rm Mpc}$ unit.
To perform the MCMC up to the smallest possible radius allowed by ICs resolution, we use the numerical solution of the first crossing problem, Eqs. (3.4)–(3.5), implemented with the 3-parameters barrier Eq. (4.1). To obtain a numerical multiplicity function with enough precision to be used in such MCMC, we use $10^{8}$ MC realisations of Eq. (3.4) for each $\alpha$, $\beta$, and $\gamma$ values, computed with a binning radius size at least 4 times smaller than the binning used for the VFS measured in simulations ICs. This guarantees a numerical precision that is, at worst, less than $1/10$ the Poissonian uncertainty of the measured VSF. Since this is computationally expensive for a MCMC analysis, we interpolate the numerical multiplicity function over a sample of $\alpha$, $\beta$, and $\gamma$ values. The computational cost scales linearly with the number of realisations performed and quadratically with the number of radius bins considered. For reference, with the above specifications, the first crossing distribution computed with $10^{8}$ realisations in a random position in the parameter space within domino for the $\delta_{\rm v}=-0.62$ case, and 109 radius bins, takes around 13 seconds on a core with 48 CPUs at 2.90GHz. To select the range in a convenient way, we produce a large number of numerical multiplicity over a random sapling of $\alpha$, $\beta$, and $\gamma$, selected within the parameter space region delimited by the MCMC analysis involving the analytical multiplicity function, as described before.
At each MCMC step, we interpolate over this sample finding the closest tetrad in which the point of coordinates ($\alpha$, $\beta$, $\gamma$) is embedded (which in most of the cases does not correspond to the 4 closest sample points). For interpolating, we weight the numerical multiplicity function corresponding to each point using the barycenter coordinates. The likelihood is the same as Eq. (E.2), however, the theoretical model is now obtained by numerically integrating over the 4 theoretical radius bins composing one data radius bin, i.e. $n^{T}_{i}(\Theta)=\sum_{j=1}^{4}N_{\times}(R_{j})/N_{\rm tot}$, with $N_{\rm tot}=10^{8}$, where $N_{\times}(R_{j})$ is the number of crossing interpolated with the barycenter coordinates. In addition, the variance is now the square sum of the numerical uncertainty of the theoretical model and the Poissonian uncertainty of measurements from simulations: $\Sigma_{i}=N_{i}\left[4\pi(\langle R\rangle_{i}/L_{\rm box})^{3}/3\right]^{2}+\sum_{j=1}^{4}N_{\times}(R_{j})/N_{\rm tot}^{2}$. We verified that, for this specific application, this interpolation method is more accurate than the standard ones over a grid. Moreover, this approach allows to simply increase the sample resolution if necessary, by producing more realisations at random positions within the same parameter space region, without worrying about the grid.
As a last consideration, it is worth noting that the moving barrier shape, Eq. 4.1, is robust in modelling the Lagrangian VSF measured from simulation ICs. In particular, the five larger radius bins are always enough to constrain the 3-parameter barrier and to predict the VSF down to the smaller radii. However, using all the available radius bins, the uncertainty in the posterior decreases.
All the MCMC analyses are performed using the emcee sampler [98].
Appendix F Redshift dependence of the HMF
@@
Figure 9: Dependence of the halo formation threshold $\delta_{\rm h}$ and the moving barrier parameter parameters of Eq. (4.1) as a function of $\sigma_{8}(z)$. The upper axis shows the corresponding redshift. Orange violins show the posterior distribution obtained at each DEMNUni snapshot considered, $z=0$, 0.47, 1.05, 1.46, 2.05. Blue solid lines and dashed areas show the best-fit and 68% CL of the scaling relations described in the text.
@@
In this Section we discuss the redshift dependence of the Eulerian HMF, as described in Sec. 5, obtained using Eq. (4.2) as prior to fit the formation barrier $\delta_{\rm h}$. The redshift dependence of the theoretical HMF is better expressed as a function of $\sigma_{8}(z)=\sqrt{C(R,R)}$ at a given redshift, with $R=8\,h^{-1}{\rm Mpc}$. In particular, $\sigma_{8}(z)=\sigma_{8}(z=0)D_{0}(z)/D_{0}(z=0)$, where $D_{0}(z)$ is the linear growth factor of perturbations. In this way, the normalisation effects of the Lagrangian space and any redshift dependence are automatically included. Nevertheless, Fig. 9 shows that all the parameters of the effective barrier, Eq. (4.1) and $\delta_{\rm h}$ can be consistently considered redshift independent. The orange violins shows the posterior of the parameter at the $\sigma_{8}(z)$ values corresponding to the redshift values explored, $z=0$, 0.47, 1.05, 1.46, 2.05. Blue lines show the best-fit value and the shaded area the 68% CL. The best-fit values are the following:
$\delta_{\rm h}$ $=2.424\pm^{0.076}_{0.071}$ $\alpha$ $=1.209\pm^{0.008}_{0.007}$ (F.1)
@@
It is worth noting that redshift independent barrier parameters entail a self-similar HMF, i.e. an HMF that depends on $S=\langle\delta^{2}\rangle$ (and $D$) values only. Note that the Lagrangian space is the initial density field linearly evolved up to the redshift of interest with the linear growth factor. This means that the linear formation threshold $\delta_{\rm h}$ does not depend on the redshift, but the linear growth factor used to evolve the Lagrangian space does. In computations, this can be considered alternatively by using the linear power spectrum of the redshift considered, or using the power spectrum at a reference redshift, e.g. $z=0$, and extrapolating the formation threshold at the redshift of interest: $\delta_{\rm h}\rightarrow\delta_{\rm h}D_{0}(z=0)/D_{0}(z)$  [SMT]. According to the same argument, the effective moving barrier can be used either with the linear power spectrum at the redshift considered, or with the power spectrum at a reference redshift and extrapolating the barrier, which means $\alpha\rightarrow\alpha D_{0}(z=0)/D_{0}(z)$, where we considered $z=0$ as the reference redshift. Moreover, any parameter multiplied or divided by $\sigma$, or more generally by any quantity derived from the power spectrum, must be corrected for the growth factor. Therefore, also $\beta\rightarrow\beta D_{0}(z=0)/D_{0}(z)$.
@@
Figure 10: Relative difference between theoretical HMFs and with respect to the HMF measured in the DEMNUni in unit of Poissonian uncertainty: Eulerian spherical mapping of the effective barrier approach implemented in this work (blue), Eqs. (3.2), (4.1), fitting the halo threshold from Eq. (4.2); SMT best-fit model (orange); Tinker et al. 2008 [63] best-fit model (green). Each panel shows a different redshift, $z=0$, 0.47, 1.05, 1.46, 2.05. Errorbars and hatched area show ±$\pm 1\sigma$ interval, dashed area shows the ±$\pm 1\%$ uncertainty.
The blue curve in Fig. 10 shows the relative difference between the Eulerian spherical mapping of the effective barrier approach implemented in this work (blue), Eqs. (3.2), (4.1), fitting the halo threshold from Eq. (4.2) as described in Sec. 5. Each panel shows a different redshift, corresponding to the DEMNUni snapshots considered. The error bars and the hatched area show ±$\pm 1\sigma$ interval, dashed area shows the ±$\pm 1\%$ uncertainty. The agreement is almost always within $1\sigma$. Similarly to the case of fitting the moving barrier for the Lagrangian VSF, Appendix E, the model presented here is robust in modelling the Eulerian HMF measured from simulation. As in the previous case, the posterior distributions of the considered parameter are stable from the larger scale explored down to the resolution limit of the simulation. For comparison, the best-fit of the SMT model (orange lines) and the best-fit of the 4-parameters Tinker et al. 2008 [63] model are shown. It can be noticed that the agreement of our model with the HMF measured in simulation is almost always as good as the Tinker et al. 2008 [63] one. The main differences are the clearer physical interpretation and the smaller parameter space of the model presented in this work. Moreover, our model converges much faster in the MCMC fit than the 4-parameters Tinker et al. 2008 [63] model. Note that the multiplicity functions of Eq. (3.14) and Eq. (3.2) can be also used as a versatile fitting function with as many fitting parameters as desired, which modify $B_{s}$ and $B^{\prime}_{S}$ quantities.